# Federated Neural Graph Databases

### v3

@@ 3.1. Preliminary
3.2. Complex Logical Query
The complex logical query is defined in existential positive first-order logic form, consisting of various types of logic expressions like existential quantifiers $\exists$, logic conjunctions $\wedge$, and disjunctions $\vee$. In the logical expression, there is a set of anchor entities $V_{a}\in\mathcal{V}$ denotes given context, existential quantified variables $V_{1},V_{2},...,V_{k}\in\mathcal{V}$, and a unique variable $V_{?}$ denotes our query target. The complex query intends to find the target answers $V_{?}\in\mathcal{V}$, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the graph-structured data that can satisfy the given logical expressions simultaneously. Following the definition in (Ren et al., 2020), the complex query expression can be converted to the disjunctive normal form (DNF) in the following:
(1) $\begin{split}q[V_{?}]&=V_{?}.V_{1},...,V_{k}:c_{1}\lor c_{2}\lor...\lor c_{n}\\ c_{i}&=e_{i,1}\land e_{i,2}\land...\land e_{i,m},\end{split}$
@@ 3.2. Complex Logical Query
Figure 2. The training and retrieval process of FedNGDB. The blue line denotes the training process, and the green line denotes the retrieval process. In the training, clients’ NGDB models are trained on respective graph-structured data. At each round, local NGDBs are aggregated at the server and updated using the global parameters. Among them, the embeddings are protected by secret aggregation so that the server can not access them. In the retrieval, each query is decomposed into sub-queries. Clients compute sub-query embeddings which the server is used to aggregate query embeddings. Answer scores are computed at clients and are aggregated at a server to retrieve answers.
3.3. Distributed Graph Set Query
@@
Definition 3.0 (Cross-graph Query).
@@
For example in figure 1, the entity ”Physics” is the answer to the query $q$ because there exists an existentially quantified variable ”Einstein” that can satisfy the logical expression. The query $q$ is a cross-graph query as the atomic expressions in the query are from different graph databases and can not be found in a single graph. For example, Win(Einstein, Nobel Prize) and BornIn(Einstein, Germany) are two atomic expressions from different graph databases. When the distributed graph databases face cross-graph queries, the answer can not be inferred from a single graph. Besides, We have the definition of in-graph query correspondingly:
Definition 3.0 (In-graph Query).
@@ 3.4. Problem Formulation
4. Federated Neural Graph Databases
@@ 4.1. Model Learning
4.1.1. Secret Aggregation
There are various techniques, like homomorphic encryption (HE) (Paillier, 1999), secure multi-party computation (MPC) (Mohassel and Zhang, 2017), and differential privacy (Dwork, 2008; Geyer et al., 2017) to protect the uploaded parameters, however the protection of aggregated global model are often ignored. Unfortunately, the global model is informative and vulnerable to privacy attacks  (Zhang et al., 2022; Hu and Song, 2023). Hence we propose a secret aggregation applied in parameter aggregation which can prevent the global server from knowing the aggregated parameters using homomorphic encryption. Assume that at each client $i$, a parameter denoted as $\theta_{i}$ is uploaded to a server for aggregating global parameter $\theta$. The procedure of secret aggregation is described in Algorithm 1. In the beginning, each client $C_{i}$ randomly generates perturbed parameters $\theta_{i}^{r}$ and shares them with other clients with encryption (Shown in Appendix A). After the sharing, each client has a set of parameters $\{\theta_{1}^{r},\theta_{2}^{r}.\cdots,\theta_{n}^{r}\}$. In the training process, for each client, $C_{i}$ uploads perturbed parameter $(\theta_{i}+\theta_{i}^{r})$ to the server under the protection of homomorphic encryption. The server collects all perturbed parameters from clients, computes aggregated perturbed parameter $\theta^{r}$, and sends it back to all clients. Finally, after receiving the perturbed parameters, the clients first decrypt the parameters and can compute aggregated parameter $\theta$ by removing the perturbed parameters and preventing exposing it to the server.
4.1.2. Model Training
Similar to (Chen et al., 2021), the server constructs a set of mapping matrices $\{\mathbf{M}^{i}\ \in\{0,1\}^{n\times n_{i}}\}_{i=1}^{N}$ and existence vectors $\{\mathbf{v}^{i}\in\{0,1\}^{n\times 1}\}_{i=1}^{N}$ to denote the entities in each client, where $n$ is the number of all unique entities in KG set and $n_{i}$ is the number of entities from client $C_{i}$. $\mathbf{M}^{i}_{m,n}=1$ if the $m$-th entity in entity table corresponds to the $n$-th entity from client $C_{i}$. $\mathbf{v}^{i}_{m}=1$ indicates that the $m$-th entity in entity table exists in client $C_{i}$.
@@ 4.1.2. Model Training
(2) $\mathbf{E}_{t+1}^{r}\leftarrow\left(\mathbbm{1}\oslash\sum_{i\in\mathbf{C}}\mathbf{v}^{i}\right)\otimes\sum_{i\in\mathbf{C}}\mathbf{M}^{i}(\mathbf{E}_{t+1}^{i}+\mathbf{E}_{r}^{i}),$
@@
(3) $\mathbf{E}_{t+1}^{r,i}\leftarrow{\mathbf{M}^{i}}^{\top}\mathbf{E}_{t+1}^{r},$
@@
(4) $\mathbf{E}_{t+1}^{i}\leftarrow\mathbf{E}_{t+1}^{r,i}-{\mathbf{M}^{i}}^{\top}\left(\mathbbm{1}\oslash\sum_{j\in\mathbf{C}}\mathbf{v}^{j}\right)\otimes\sum_{j\in\mathbf{C}}\mathbf{M}^{j}\mathbf{E}^{j}_{r}.$
@@ 4.2.2. Answer Retrieval
(5) $\mathbf{S}^{i}_{q}\leftarrow f^{i}_{s}(\mathcal{V}_{i},q)\in\mathbb{R}^{n_{i}\times 1},$
@@
(6) $\mathbf{S}\leftarrow\left(\mathbbm{1}\oslash\sum_{i=1}^{N}\mathbf{v}^{i}\right)\otimes\sum_{i=1}^{N}\mathbf{M}^{i}\mathbf{S}^{i}.$
@@ 5. Experiments
5.1. Datasets and Experiment Setting
@@
Figure 3. The query structures used for evaluation in the experiments. Naming for each query structure is provided under each subfigure, for brevity, the p, i, and u represent the projection, intersection, and union operations respectively.
5.1.1. Datasets
In our experiment, following previous work, we use the three commonly used knowledge graphs as graph-structured data: FB15k (Bollacker et al., 2008; Bordes et al., 2013), FB15k-237 (Toutanova and Chen, 2015), and NELL995 (Carlson et al., 2010) to construct the distributed query answering benchmark. In each dataset, there are vertices describing entities and edges describing relations. To evaluate the distributed complex query, we conduct experiments assuming having 3 and 5 clients in each federated neural graph database system, respectively. We randomly select relations into clients and split triples into clients according to selected relations as a local graph database. The triplets in each local graph are separated into training, validation, and testing with a ratio of 8:1:1 respectively. Following previous works (Ren et al., 2020), we construct training graph $\mathcal{G}_{train}$, validation graph $\mathcal{G}_{val}$, and test graph $\mathcal{G}_{test}$ in each client by training edges, training+validation edges, and training+validation+testing edges, respectively. The detailed statistics are listed in the table 1. #Clients denotes the number of clients, #Nodes, #Relations, #Edges denote the average number of nodes, relations, and edges in each client respectively.
Table 1. The statistics of three datasets used for experiments.
@@ 5.1.1. Datasets
5.1.2. Query Sampling
Following previous work (Hamilton et al., 2018; Bai et al., 2023b), we evaluate the complex logical query answering performance on the following eight general query types with abbreviations of $1p,2p,2i,ip,$ $3i,pi,2u$, and $up$. As shown in figure 3, each subgraph denotes one query type, where each edge represents either a projection or a logical operator, and each node represents either a set of entities, the anchor entities and relations are to be specified to instantiate logical queries. We use the sampling method commonly used in previous works (Bai et al., 2023c; Ren et al., 2020) to randomly sample complex queries from graphs. We randomly sample two sets of queries from the graph sets: in-graph queries and cross-graph queries to evaluate local and global answer retrieval performance. For local model evaluation, we first obtain training, validation, and testing queries from the formerly constructed local graph databases respectively. Then for the training queries, we conduct a graph search to find corresponding training answers on the local training graph. For the validation queries, we search for the answers on both the training graph and the validation graph and only use those queries that have different numbers of answers on the two graphs. For the testing queries, we use those queries that have different answers on the testing graph from answers on the training graph and validation graph. For global model evaluation, we construct global training, validation, and testing graphs using all local graphs, and sample testing queries with atomic expressions from different local graphs, finally we search for answers on three global graphs and only use those queries that have different answers on the testing graph from other two graphs. We collect statistics of complex queries in three datasets and the statistics are shown in Table 2. The number of in-graph queries is the average number of the client’s local queries.
Table 2. The statistics of queries sampled from three datasets used for experiments.
@@ 5.1.2. Query Sampling
Table 3. The retrieval performance of distributed neural graph databases when there are 3 clients. The average results of HR@3 and MRR of all clients are reported. The best results are underlined. The best results of distributed models are in bold.
@@
[…]
5.1.3. Baselines
We can use various existing query encoding methods as our local base model, to evaluate the effectiveness and generalization ability of our proposed FedNGDB, we select three commonly used complex encoding methods GQE (Hamilton et al., 2018), Q2P (Bai et al., 2022), Tree-LSTM (Bai et al., 2023c) as our base model. GQE is a graph query encoding model that encodes a complex query into a vector in embedding space; Q2P represents complex queries using multiple vectors; Tree-LSTM recursively represents complex queries and treats all operations, entities, and relations as tokens.
To the best of our knowledge, there are no existing federated complex query answering methods but several federated knowledge graph embedding methods, therefore, we choose to compare our methods with FedE (Chen et al., 2021) and FedR (Zhang et al., 2022), two commonly used federated knowledge graph embedding methods as baselines. FedE aggregates both entity embeddings and relation embeddings in a server, while FedR only aggregates relation embeddings for privacy concerns and communication efficiency. We utilize these two methods with slight modifications to train a global complex query answering model: FedE aggregates all query encoding parameters and FedR aggregates relation embeddings and query encoding networks. Besides, we also compare our FedNGDB with local and central settings. In the local setting, there is no collaboration between clients, while in the central setting, all distributed graphs in the graph set are aggregated for a global graph for training, we sample complex queries from global training, and validation graphs for training and validation.
If there is no further statement, we use the following implementation settings in the experiments. We tune hyper-parameters on the validation local queries for the base query encoding methods and set the dimension of entities and relations as 400 for all models for fair comparison and use AdamW (Loshchilov and Hutter, 2018) as optimizer.
Table 4. The retrieval performance of distributed neural graph databases when there are 5 clients. The average results of HR@3 and MRR of all clients are reported. The best results are in bold.
FB15k-237 FB15k NELL995
@@ 5.1.3. Baselines
Tree-LSTM FedNGDB 13.79 13.27 12.74 12.18 15.44 15.81 15.28 14.24 15.68 14.28 14.57 12.89
5.1.4. Evaluation Metrics
@@
(7) $\text{Metric}(q)=\frac{1}{|M_{test}/M_{val}|}\sum_{v\in M_{test}/M_{val}}m(rank(v)),\$
@@
(a) Cross-graph Queries
(b) Cross-graph Queries
(c) In-graph Queries
(d) In-graph Queries
Figure 4. The evaluation results of FedNGDB-GQE facing different types of cross-graph queries on FB15k-237 (subfigures (a), (b)). The evaluation results of FedNGDB-GQE model facing in-graph queries on FB15k-237 (subfigures (c), (d)).
5.2. Performance Evaluation
@@
5.3. Query Types
FedNGDB can globally reason over distributed graphs and retrieve answers to cross-graph queries. To evaluate FedNGDB’s performance on various types of complex queries, we conduct experiments to evaluate the retrieval performance of FedNGDB and compare it to central learning on FB15k-237 when there are 3 clients. Figure 4(a) and Figure 4(b) show the FedNGDB-GQE’s performance facing cross-graph queries. As we can see from the figure, FedNGDB performs well on most various types of queries compared to the central model. For example, on query types ’2i’ and ’pi’, FedNGDB can reach more than 90% MRR compared to the central model.
5.4. Local Influence
Because the in-graph queries can be processed by a single local neural graph database, in this part, we evaluate the performance of FedNGDB on these queries to assess the influence of FedNGDB on local queries. We conduct experiments on FB15k-237 and the number of clients is 5. We evaluate the performance of FedNGDB based on GQE and compare the model with no collaboration. The results are summarized in the Figure 4(c) and Figure 4(d). As shown in the figure, although each client has different performance due to the triplets correlation being different in each sub-dataset, FedNGDB can improve all clients’ performance compared to the model without collaboration.
5.5. Convergence Rate
We evaluate the convergence speed of three federated frameworks. The results are presented by the average number of communication round ratios relative to FedE. As shown in Table 5, FedC’s convergence speed is faster than FedC while slightly slower than FedE, demonstrating that our FedNGDB can protect stronger protection while remaining competitive efficiency.
Table 5. The statistics of communication rounds .
Setting FedE FedR FedNGDB
Relative Rounds to FedE 1.00 1.32 1.09
6. Conclusion

### v4

@@ 3.1 Preliminary
3.2 Complex Logical Query
The complex logical query is defined in existential positive first-order logic form, consisting of various types of logic expressions like existential quantifiers $\exists$, logic conjunctions $\wedge$, and disjunctions $\vee$. In the logical expression, there is a set of anchor entities $V_{a}\in\mathcal{V}$ denotes given context, existential quantified variables $V_{1},V_{2},...,V_{k}\in\mathcal{V}$, and a unique variable $V_{?}$ denotes our query target. The complex query intends to find the target answers $V_{?}\in\mathcal{V}$, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the graph-structured data that can satisfy the given logical expressions simultaneously. Following the definition in Ren et al. (2020), the complex query expression can be converted to the disjunctive normal form (DNF) in the following:
$\begin{split}q[V_{?}]&=V_{?}.V_{1},...,V_{k}:c_{1}\lor c_{2}\lor...\lor c_{n}\\ c_{i}&=e_{i,1}\land e_{i,2}\land...\land e_{i,m},\end{split}$ (1)
@@ 3.2 Complex Logical Query
Figure 2: The training and retrieval process of FedNGDB. The blue line denotes the training process, and the green line denotes the retrieval process. In the training, clients’ NGDB models are trained on respective graph-structured data. At each round, local NGDBs are aggregated at the server and updated using the global parameters. Among them, the embeddings are protected by secret aggregation so that the server can not access them. In the retrieval, each query is decomposed into sub-queries. Clients compute sub-query embeddings which the server is used to aggregate query embeddings. Answer scores are computed at clients and are aggregated at a server to retrieve answers.
3.3 Distributed Graph Set Query
@@
Definition 3.1 (Cross-graph Query)
@@
For example in figure 1, the entity "Physics" is the answer to the query $q$ because there exists an existentially quantified variable "Einstein" that can satisfy the logical expression. The query $q$ is a cross-graph query as the atomic expressions in the query are from different graph databases and can not be found in a single graph. For example, Win(Einstein, Nobel Prize) and BornIn(Einstein, Germany) are two atomic expressions from different graph databases. When the distributed graph databases face cross-graph queries, the answer can not be inferred from a single graph. Besides, We have the definition of in-graph query correspondingly:
Definition 3.2 (In-graph Query)
@@ 3.4 Problem Formulation
4 Federated Neural Graph Databases
@@ 4.1 Model Learning
4.1.1 Secret Aggregation
There are various techniques, like homomorphic encryption (HE) (Paillier, 1999), secure multi-party computation (MPC) (Mohassel & Zhang, 2017), and differential privacy (Dwork, 2008; Geyer et al., 2017) to protect the uploaded parameters, however the protection of aggregated global model are often ignored. Unfortunately, the global model is informative and vulnerable to privacy attacks  (Zhang et al., 2022; Hu & Song, 2024). Hence we propose a secret aggregation applied in parameter aggregation which can prevent the global server from knowing the aggregated parameters using homomorphic encryption. Assume that at each client $i$, a parameter denoted as $\theta_{i}$ is uploaded to a server for aggregating global parameter $\theta$. The procedure of secret aggregation is described in Algorithm 1. In the beginning, each client $C_{i}$ randomly generates perturbed parameters $\theta_{i}^{r}$ and shares them with other clients with encryption (Shown in Appendix A). After the sharing, each client has a set of parameters $\{\theta_{1}^{r},\theta_{2}^{r}.\cdots,\theta_{n}^{r}\}$. In the training process, for each client, $C_{i}$ uploads perturbed parameter $(\theta_{i}+\theta_{i}^{r})$ to the server under the protection of homomorphic encryption. The server collects all perturbed parameters from clients, computes aggregated perturbed parameter $\theta^{r}$, and sends it back to all clients. Finally, after receiving the perturbed parameters, the clients first decrypt the parameters and can compute aggregated parameter $\theta$ by removing the perturbed parameters and preventing exposing it to the server.
Besides, the parameters are also protected by the differential privacy to prevent the potential privacy leakage (Zhu et al., 2019), we use local differential privacy (LDP) to the local gradients to further improve the difficulty to recover sensitive information from transmission. As proposed in Yi et al. (2021), we clip the local gradients based on the $L_{\infty}$ norm with the threshold $\mathcal{C}$ and apply zero-mean Laplacian noise with strength $\lambda$, the upper bound of the privacy budget $\epsilon$ is $\frac{2\mathcal{C}}{\lambda}$. The detailed discussion can be found in Appendix B.3.
4.1.2 Model Training
Similar to Chen et al. (2021), the server constructs a set of mapping matrices $\{\mathbf{M}^{i}\ \in\{0,1\}^{n\times n_{i}}\}_{i=1}^{N}$ and existence vectors $\{\mathbf{v}^{i}\in\{0,1\}^{n\times 1}\}_{i=1}^{N}$ to denote the entities in each client, where $n$ is the number of all unique entities in KG set and $n_{i}$ is the number of entities from client $C_{i}$. $\mathbf{M}^{i}_{m,n}=1$ if the $m$-th entity in entity table corresponds to the $n$-th entity from client $C_{i}$. $\mathbf{v}^{i}_{m}=1$ indicates that the $m$-th entity in entity table exists in client $C_{i}$.
@@ 4.1.2 Model Training
$\mathbf{E}_{t+1}^{r}\leftarrow\left(\mathbbm{1}\oslash\sum_{i\in\mathbf{C}}\mathbf{v}^{i}\right)\otimes\sum_{i\in\mathbf{C}}\mathbf{M}^{i}(\mathbf{E}_{t+1}^{i}+\mathbf{E}_{r}^{i}),$ (2)
@@
$\mathbf{E}_{t+1}^{r,i}\leftarrow{\mathbf{M}^{i}}^{\top}\mathbf{E}_{t+1}^{r},$ (3)
@@
$\mathbf{E}_{t+1}^{i}\leftarrow\mathbf{E}_{t+1}^{r,i}-{\mathbf{M}^{i}}^{\top}\left(\mathbbm{1}\oslash\sum_{j\in\mathbf{C}}\mathbf{v}^{j}\right)\otimes\sum_{j\in\mathbf{C}}\mathbf{M}^{j}\mathbf{E}^{j}_{r}.$ (4)
@@ 4.2.2 Answer Retrieval
$\mathbf{S}^{i}_{q}\leftarrow f^{i}_{s}(\mathcal{V}_{i},q)\in\mathbb{R}^{n_{i}\times 1},$ (5)
@@
$\mathbf{S}\leftarrow\left(\mathbbm{1}\oslash\sum_{i=1}^{N}\mathbf{v}^{i}\right)\otimes\sum_{i=1}^{N}\mathbf{M}^{i}\mathbf{S}^{i}.$ (6)
@@ 5 Experiments
5.1 Datasets and Experiment Setting
@@
Figure 3: The query structures used for evaluation in the experiments. Naming for each query structure is provided under each subfigure, for brevity, the p, i, and u represent the projection, intersection, and union operations respectively.
5.1.1 Datasets
In our experiment, following previous work, we use the three commonly used knowledge graphs in central graph reasoning tasks as graph-structured data: FB15k (Bollacker et al., 2008; Bordes et al., 2013), FB15k-237 (Toutanova & Chen, 2015), and NELL995 (Carlson et al., 2010) to construct the distributed query answering benchmark, which can make better comparison to existing works. In each dataset, there are vertices describing entities and edges describing relations. To evaluate the distributed complex query, we follow the common settings of federated knowledge graph emebddings (Chen et al., 2021; Zhang et al., 2022), conducting experiments assuming having 3 and 5 clients in each federated neural graph database system, respectively. For more clients, we also conduct simple exeperiments as shown in Appendix C.1. We randomly select relations into clients and split triples into clients according to selected relations as a local graph database. The triplets in each local graph are separated into training, validation, and testing with a ratio of 8:1:1 respectively. Following previous works (Ren et al., 2020), we construct training graph $\mathcal{G}_{train}$, validation graph $\mathcal{G}_{val}$, and test graph $\mathcal{G}_{test}$ in each client by training edges, training+validation edges, and training+validation+testing edges, respectively. The detailed statistics are listed in the table 1. #Clients denotes the number of clients, #Nodes, #Relations, #Edges denote the average number of nodes, relations, and edges in each client respectively.
Table 1: The statistics of three datasets used for experiments.
@@ 5.1.1 Datasets
5.1.2 Query Sampling
Following previous work (Hamilton et al., 2018; Bai et al., 2023b), we evaluate the complex logical query answering performance on the following eight general query types with abbreviations of $1p,2p,2i,ip,$ $3i,pi,2u$, and $up$. As shown in figure 3, each subgraph denotes one query type, where each edge represents either a projection or a logical operator, and each node represents either a set of entities, the anchor entities and relations are to be specified to instantiate logical queries. We use the sampling method commonly used in previous works Bai et al. ; Ren et al. (2020) to randomly sample complex queries from graphs. We randomly sample two sets of queries from the graph sets: in-graph queries and cross-graph queries to evaluate local and global answer retrieval performance. For local model evaluation, we first obtain training, validation, and testing queries from the formerly constructed local graph databases respectively. Then for the training queries, we conduct a graph search to find corresponding training answers on the local training graph. For the validation queries, we search for the answers on both the training graph and the validation graph and only use those queries that have different numbers of answers on the two graphs. For the testing queries, we use those queries that have different answers on the testing graph from answers on the training graph and validation graph. For global model evaluation, we construct global training, validation, and testing graphs using all local graphs, and sample testing queries with atomic expressions from different local graphs, finally we search for answers on three global graphs and only use those queries that have different answers on the testing graph from other two graphs. We collect statistics of complex queries in three datasets and the statistics are shown in Table 2. The number of in-graph queries is the average number of the client’s local queries.
Table 2: The statistics of queries sampled from three datasets used for experiments.
@@ 5.1.2 Query Sampling
Table 3: The retrieval performance of distributed neural graph databases when there are 3 clients. The average results of HR@3 and MRR of all clients are reported. The best results are underlined. The best results of distributed models are in bold.
@@
NELL995 FedNGDB 14.21 13.27 13.76 12.67 16.62 15.28 16.27 16.23 16.28 15.38 16.09 15.27
5.1.3 Baselines
We can use various existing query encoding methods as our local base model, to evaluate the effectiveness and generalization ability of our proposed FedNGDB, we select three commonly used complex encoding methods GQE (Hamilton et al., 2018), Q2P (Bai et al., 2022), Tree-LSTM (Bai et al., ) as our base model. GQE is a graph query encoding model that encodes a complex query into a vector in embedding space; Q2P represents complex queries using multiple vectors; Tree-LSTM recursively represents complex queries and treats all operations, entities, and relations as tokens.
To the best of our knowledge, there are no existing federated complex query answering methods but several federated knowledge graph embedding methods, therefore, we choose to compare our methods with FedE (Chen et al., 2021) and FedR (Zhang et al., 2022), two commonly used federated knowledge graph embedding methods as baselines. FedE aggregates both entity embeddings and relation embeddings in a server, while FedR only aggregates relation embeddings for privacy concerns and communication efficiency. We utilize these two methods with slight modifications to train a global complex query answering model: FedE aggregates all query encoding parameters and FedR aggregates relation embeddings and query encoding networks. Besides, we also compare our FedNGDB with local and central settings. These two baselines represent the lower bound and upper bound, respectively, in terms of the amount of information available to the training system. In the local setting, there is no collaboration between clients, while in the central setting, all distributed graphs in the graph set are aggregated for a global graph for training, we sample complex queries from global training, and validation graphs for training and validation.
If there is no further statement, we use the following implementation settings in the experiments. We tune hyper-parameters on the validation local queries for the base query encoding methods and set the dimension of entities and relations as 400 for all models for fair comparison and use AdamW (Loshchilov & Hutter, 2018) as optimizer. We set the gradient clip threshold $C=0.1$ and Laplacian noise scale $\lambda=0.2$ to achieve $1$-DP.
Table 4: The retrieval performance of distributed neural graph databases when there are 5 clients. The average results of HR@3 and MRR of all clients are reported. The best results are underlined. The best results of distributed models are in bold.
GQE Q2P Tree-LSTM
@@ 5.1.3 Baselines
Central 14.56 14.28 13.75 13.42 15.63 15.14 15.27 14.75 16.47 15.92 15.74 14.47
NELL995 FedNGDB 13.79 13.27 12.74 12.18 15.44 15.81 15.28 14.24 15.68 14.28 14.57 12.89
5.1.4 Evaluation Metrics
@@
$\text{Metric}(q)=\frac{1}{|M_{test}/M_{val}|}\sum_{v\in M_{test}/M_{val}}m(rank(v)),\$ (7)
@@
5.2 Performance Evaluation
@@
5.3 Query Types
FedNGDB can globally reason over distributed graphs and retrieve answers to cross-graph queries. To evaluate FedNGDB’s performance on various types of complex queries, we conduct experiments to evaluate the retrieval performance of FedNGDB and compare it to central learning on FB15k-237 when there are 3 clients. Figure 4(a) shows the FedNGDB-GQE’s performance facing cross-graph queries. As we can see from the figure, FedNGDB performs well on most various types of queries compared to the central model. For example, on query types ’2i’ and ’pi’, FedNGDB can reach more than 90% MRR compared to the central model.
(a) Cross-graph Queries
(b) In-graph Queries
Figure 4: The evaluation results of FedNGDB-GQE facing cross-graph queries on FB15k-237 (subfigures (a)). The evaluation results of FedNGDB-GQE facing in-graph queries on FB15k-237 (subfigures (c)).
5.4 Local Influence
Because the in-graph queries can be processed by a single local neural graph database, in this part, we evaluate the performance of FedNGDB on these queries to assess the influence of FedNGDB on local queries. We conduct experiments on FB15k-237 and the number of clients is 5. We evaluate the performance of FedNGDB based on GQE and compare the model with no collaboration. The results are summarized in the Figure 4(b). As shown in the figure, although each client has different performance due to the triplets correlation being different in each sub-dataset, FedNGDB can improve all clients’ performance compared to the model without collaboration.
6 Conclusion